OpenStudy (anonymous):

Need help with 8 questions please

OpenStudy (anonymous):

Which equation represents a nonlinear function?

OpenStudy (anonymous):

answers

OpenStudy (anonymous):

remember: a line is represented by: y = mx + b where m is the slope and b is the y-intercept.

OpenStudy (anonymous):

so the answer would be B

OpenStudy (anonymous):

\[y = 3x^2 + 1 \] this one is ^2 so it would be a parabola not a line.

OpenStudy (anonymous):

hence: nonlinear.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

yep the third one.

OpenStudy (anonymous):

question

OpenStudy (anonymous):

answers

OpenStudy (anonymous):



OpenStudy (anonymous):

so d

OpenStudy (anonymous):

In general, a linear function is a function defined by an equation of the form f(x) = y= mx+b and the graph of a linear equation is a straight line. In the above equation, "m" is called the slope of the line and "b" is called the intercept on y-axis.

OpenStudy (anonymous):

TRUE: --x increases by 1 each time. --y increases by a different value each time. -- the rate of change is not constant FALSE: It shows a linear function.

OpenStudy (anonymous):

@some_someone sorry to disagree but a linear function has the form y=mx. The form y=mx+b is a line but is not linear. Linear functions y=f(x) fulfill: \[f(m·x_1+n·x_2)=mf(x_1)+nf(x_2)\] According to this defintion, none of the functions showed is a linear function. A function y=mx+b is called "affine function" and is non-linear

OpenStudy (anonymous):

Hmmm. okaii so based on your thoughts, what is the answer? hmmmm?

OpenStudy (anonymous):

they are not my thoughts, they are formal definitions. None is linear. The problem seems to be aiming at the quadratic function (C). If this is the case I would dare say there is a conceptual error in the teacher who has defined the exercise, when considering y=mx+b is linear. Simply double the x and see if the y doubles too and it does not

OpenStudy (anonymous):

If they are asking whether it is a line or not, that is a horse of a different color. But being a line does not mean is linear.